7.10. Steam expands adiabatically in a nozzle from inlet conditions of 130(psia), 420(°F), 
and a velocity of 230(ft)(s)−1 to a discharge pressure of 35(psia) where its velocity is 
2000(ft)(s)−1. What is the state of the steam at the nozzle exit, and what is Sg for the 
process?
Answer #1

inlet temperature T1 = 420°F

Inlet pressure P1 = 130 psia

velocity of inlet steam u1 =

Velocity of outlet steam u2 = 2000 ft/s

From the steam table

From the energy balance

For an adiabatic process

For horizontal position

final energy balance

H2 - H1 = (230² - 2000²)/2 = - 1973550 ft2/s2 x (1Btu/lbm / 25.037 ft2/s2)

H2 = 1233.6 - 78.825

H2 = 1154.775 Btu/lbm

At 35 psi

From the steam table

Hf = 228.03 Btu/lbm

Hg = 1167.1 Btu/lbm

Sf = 0.3809 Btu/lbm-R

Sg = 1.6872 Btu/lbm-R

From the enthalpy balance

H2 = Hf + x*(Hg - Hf)

1154.775 = 228.03 + x*(1167.1 - 228.03)

x = 0.987

From the entropy balance

S2 = Sf + x*(Sg - Sf)

S2 = 0.3809 + 0.987*(1.6872 - 0.3809)

S2 = 1.67 Btu/lbm-R

Sdotg = S2 - S1

= 1.67 - 1.6310

= 0.039 Btu/lbm-R